# Stability of optimal transport on metric measure spaces

### v1

@@ 3 Stability of optimal transport maps
$\|\phi_{\mu}-\phi_{\nu}\|_{L^{\infty}(\rho)}\leq\|\phi_{\mu}\|_{L^{\infty}(\rho)}+\|\phi_{\nu}\|_{L^{\infty}(\rho)}\leq{\rm{osc}}(\phi_{\mu})+{\rm{osc}}(\phi_{\nu})\leq 2\mathop{\rm diam}\nolimits^{2}(S\cup Y).$
Combining with Theorem 1.2, we get
@@
Recall that a finite dimensional Alexandrov space is also $\rm{RCD}$ (cf. [41, 46]). Combining (3.14), Theorem 3.1 and (3.15), we obtain
@@ Appendix A Poincaré inequality: local to global
$\int_{B}|f(x)-\mathbb{E}_{\rho_{B}}(f)|\,{\mathrm{d}}\rho_{B}(x)\leq C_{4}r_{0}\int_{B}|\nabla_{x}f|\,{\mathrm{d}}\rho_{B},$ (A.1)
where $C_{4}$ depends on $K,N,a_{1},a_{2}$.
@@
By [18, Chapter 9] and [42, Remark 3.3], $\mathfrak{m}$ satisfies the $(1,1)$-Poincaré inequality:
$\int_{B}|f-\mathbb{E}_{\mathfrak{m}_{B}}(f)|\,{\mathrm{d}}\mathfrak{m}\leq c_{1}r_{0}\int_{B}|\nabla_{x}f|\,{\mathrm{d}}\mathfrak{m},$ (A.2)
@@
$\leq$ $2a_{2}c_{1}r_{0}\int_{B}|\nabla_{x}f|\,{\mathrm{d}}\mathfrak{m}\leq 2\frac{a_{2}}{a_{1}}c_{1}r_{0}\int_{B}|\nabla_{x}f|\,{\mathrm{d}}\rho,$
@@
$\int_{S}|f(x)-\mathbb{E}_{\rho}(f)|\,{\mathrm{d}}\rho(x)\leq C_{5}\sum_{B\in\mathcal{F}}\rho(B)\int_{B}|f(x)-\mathbb{E}_{\rho_{B}}(f)|\,{\mathrm{d}}\rho_{B}(x),$ (A.4)
where $C_{5}$ depends on $K,N,\mathop{\rm diam}\nolimits(S),S$.
@@
$\sum_{B\in\mathcal{F}}\rho(B)|\mathbb{E}_{\rho_{B}}(f)-\mathbb{E}_{\rho_{B_{0}}}(f)|\leq$ $2G\sum_{B\in\mathcal{F}}\rho(B)\sum_{B\subset F\overline{B}}\frac{a_{\overline{B}}}{\rho(\overline{B})}\leq 2G\sum_{\bar{B}\in\mathcal{F}}\frac{a_{\bar{B}}}{\rho(\bar{B})}\sum_{B\subset F\bar{B}}\rho(B).$ (A.9)

### v2

@@ 3 Stability of optimal transport maps
$\|\phi_{\mu}-\phi_{\nu}\|_{L^{\infty}(\rho)}\leq\|\phi_{\mu}\|_{L^{\infty}(\rho)}+\|\phi_{\nu}\|_{L^{\infty}(\rho)}\leq{\rm{osc}}(\phi_{\mu})+{\rm{osc}}(\phi_{\nu})\leq 2\big(\mathop{\rm diam}\nolimits(S\cup Y)\big)^{2}.$
Recall that a finite dimensional Alexandrov space is also $\rm{RCD}$ (cf. [41, 46]). By Theorem 1.1, we get
@@
Combining Theorem 3.1, (3.14) and (3.15), we obtain
@@ Appendix A Poincaré inequality: local to global
$\int_{B}|f(x)-\mathbb{E}_{\rho_{B}}(f)|\,{\mathrm{d}}\rho_{B}(x)\leq C_{3}r_{0}\int_{B}|\nabla f|\,{\mathrm{d}}\rho_{B},$ (A.1)
where $C_{3}$ depends on $K,N,a_{1},a_{2}$.
@@
By [18, Chapter 9] and [42, Remark 3.3], $\mathfrak{m}$ satisfies the strong local $(1,1)$-Poincaré inequality:
$\int_{B}|f-\mathbb{E}_{\mathfrak{m}_{B}}(f)|\,{\mathrm{d}}\mathfrak{m}\leq c_{1}r_{0}\int_{B}|\nabla f|\,{\mathrm{d}}\mathfrak{m},$ (A.2)
@@
$\leq$ $2a_{2}c_{1}r_{0}\int_{B}|\nabla f|\,{\mathrm{d}}\mathfrak{m}\leq 2\frac{a_{2}}{a_{1}}c_{1}r_{0}\int_{B}|\nabla f|\,{\mathrm{d}}\rho,$
@@
$\int_{S}|f(x)-\mathbb{E}_{\rho}(f)|\,{\mathrm{d}}\rho(x)\leq C_{4}\sum_{B\in\mathcal{F}}\rho(B)\int_{B}|f(x)-\mathbb{E}_{\rho_{B}}(f)|\,{\mathrm{d}}\rho_{B}(x),$ (A.4)
where $C_{4}$ depends on $K,N,\mathop{\rm diam}\nolimits(S),S$.
@@
$\sum_{B\in\mathcal{F}}\rho(B)|\mathbb{E}_{\rho_{B}}(f)-\mathbb{E}_{\rho_{B_{0}}}(f)|\leq$ $2G\sum_{B\in\mathcal{F}}\rho(B)\sum_{B\subset F\bar{B}}\frac{a_{\bar{B}}}{\rho(\bar{B})}\leq 2G\sum_{\bar{B}\in\mathcal{F}}\frac{a_{\bar{B}}}{\rho(\bar{B})}\sum_{B\subset F\bar{B}}\rho(B).$ (A.9)